The oz package^{∗}

edited by Sebastian Rahtz S.Rahtz@elsevier.co.uk

1995/03/27

Contents

1 Introduction 1 1.1 Modification History . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Introduction

This file provides macros to type Object Z schemas.

1.1 Modification History

The original zed.sty file was written by Mike Spivey. These macros have been extensively modified to include extra symbols and environments for Object-Z and now have little resemblence to the original macros. Mike Spivey has also extended his macros along different lines for Z — the latest version of macros are called fuzz.sty and come with a syntax checker for Z. They can be purchased for a small

fee from: mike@prg.oxford.ac.uk

∗This file has version number 2.4, last revised 1995/03/27.

87  original zed.sty file — Mike Spivey

88 modified to include extra symbols and environ-

ments — Paul King

88 modified to include macros for UQ editor — Ray

Neucom

May 89 modified to include object-oriented constructs —

PK

Jun 89 modified to automatically change Z symbol size

— PK

27 Jul 89 removed spurious space from \@setsize — PK
3 Aug 89 adjust style of equation number field — PK

24 Aug 89 add optional field to topline and zedline for proofs

— PK

15 Sep 89 added extra qed symbols, updated classcom and

comment — PK

25 Sep 89 renamed z@[bB]ig to z[bB]ig and changed tem-

poral symbols — PK

30 Sep 89 added \znewpage, \Infrule, removed space

above state box — PK

12 Mar 90 changed \@setsize to work better for double-

spaced text — PK

31 Mar 90 added definition for @ and \bool and redefined

‘;’ — PK

9 May 90 changed \sdef to \varsdef, \sdef & \defs to

Spivey’s latest — PK

27 May 90 added \c{n}{text} — a tab like \t{n} with text

at left — PK

29 May 90 added ‘corners’ to boxes and \zedcornerheight

— PK

11 Jul 90 added \rename*[y/x] and \zseq \zset ...,

changed \zeq \zimp — PK

2 Aug 90 added subseq — PK

9 Nov 90 changed \M to hopefully interact better with spa-

cing — PK

24 Feb 92 changed to work with NFSS — Sebastian Rahtz
29 Dec 92 changed to work with NFSS2 — Sebastian Rahtz
23 Dec 93 checked for \LaTeXe

6 Jan 94 checked again for \LaTeXe with help of David

Carlisle

7 Apr 94 checked again, bundle files added

13 Aug 94 removed hard-wired number of family for math letters removed definition of \c as it prevents cedilla from working

22 Feb 95 updated for AMSLaTeX 2.1 and Lucida compatibility

Please post any updates that you may make to this file to: Paul King – ACSnet: king@batserver.cs.uq.oz.au